for the reaction of magnesium and HCL if you react 25 grams of magnesium with 37 grams of Hcl which is reactant is the limiting reactant and how much MgCl2 will be produced by the reaction?


1
Expert's answer
2021-10-10T10:18:02-0400

"M\ng\n(\ns\n)\n+\n2\nH\nC\nl\n(\na\nq\n)\n\u2192\nM\ng\nC\nl\n2\n(\na\nq\n)\n+\nH\n2\n(\ng\n)"

Molar mass of Mg = 24.305

25/24.305= 1.029mol

Molar mass of HCl = 36.458

37/36.458= 1.015mol

Molar mass of MgCl2 = 95.211

Mole ratio of HCl: MgCl2 = 2:1 = 1.015/2

= 0.5075

Mole ratio of MgCl2:Mg= 1:1 = 1.029mol

= 1.029+0.5075= 1.5365

= 1.5365× 95.211= 146.292g